: Solve the problem and show the complete solution.

An infinite line of charge produces a field of magnitude 4.5x104 N/C at a distance of 2.0m. What is the linear charge density λ? E = λ / 2π εo r Given: field of magnitude E = 4.5x104 N/C Distance d = 2.0m εo permitivity of free space = 8.85x10-12 C2 / N.m2


1
Expert's answer
2021-02-17T11:07:58-0500

Since the electric field of an infinite charged line is given as:


"E = \dfrac{\lambda}{2\pi\varepsilon_0d}"

than the charge density is:


"\lambda = 2\pi\varepsilon_0Ed"


Substituting given values, obtain:


"\lambda = 2\pi\cdot 8.85\times 10^{-12}\cdot 4.5\times 10^4\cdot 2 \approx 5\times 10^{-6}C/m"

Answer. "5\times 10^{-6}C/m".